OpenStudy (anonymous):

A turbine at a hydroelectric power station is situated 30 m below the level of the surface of a large lake. The water passes through the turbine at a rate of 340 m3 per minute. The overall efficiency of the turbine and generator system is 90%. What is the output power of the power station? (The density of water is 1000 kg m–3.) A 0.15 MW B 1.5 MW C 1.7 MW D 90 MW

OpenStudy (anonymous):

They have given volume/time and density= so u can find the mass/time m/t = v/t* density m/t = 340/60 * 1000 then u can find change in pE/time = power input m/t * g * h = 1.67 * 10^6 the use the efficiency equation answer would be B i guess